Debts of $1,080.00 due four months ago and $1185.00 due in one month are to be settled by two equal payments due now and eight months from now respectively. Find the size of the equal payments using 12% interest and today is the focal date.
Answer #1

Rate of Interest = 12%

Focal Date = Today

Debt four months ago from today = $1080

Debt in one month from today = $1185

Lets the size of equal payment = $X

To Solve the problem bring all the cash flows at the focal date

Therefore,

1080(1.12)(4/12) + 1185/(1.12)(1/12) = X + X/(1.12)(8/12)

1121.57 + 1173.86 = X + 0.9272X

1.9272X = $2295.43

X = $1191.07

the size of equal payment = $1191.07